OpenStudy (anonymous):

How many solutions can be found for the equation 5x + 3(x − 1) = 10x − 2x − 3? None One Two Infinitely many

OpenStudy (jadeishere):

well, first you have to solve the equation 5x + 3(x - 1) = 10x - 2x - 3 5x + 3x - 3 = 10x - 2x - 3 8x - 3 = 8x - 3 0x = 0 there are infinitely many solutions because any value for the variable x would make the statement true :)